Calculate the potential for the half cell made up of zinc in 0.02M ZnSo4 solution at 298k if the So is 0.763v
1
Expert's answer
2020-11-09T13:52:37-0500

Zn2+ + 2e- = Zn

E(Zn2+/Zn) = E0(Zn2+/Zn) + R*T/(2*F)*ln([Zn2+]) =

-0.763+8.314*298/(2*96500)*ln(0.02) = -0.813 V

Answer: -0.813 V